# Energy-based generator matching: A neural sampler for general state space

### v2

@@ 1 Introduction
We tackle the problem of drawing samples from a Boltzmann distribution $p_{\text{target}}(x)\propto\exp(-\mathcal{E}(x))$ given only oracle access to the energy function $\mathcal{E}(x)$ and no pre-computed equilibrium samples. Such “energy-only” inference arises throughout machine learning, e.g., Bayesian inference [15] and statistical physics, e.g., computing thermodynamic averages [29], yet it remains intractable in high-dimensional or combinatorial state spaces. Classical methods based on Markov processes, e.g., Metropolis–Hastings [28, 18], Gibbs sampling [16], and Metropolis-adjusted Langevin algorithm [17, 33], provide asymptotically exact guarantees, but at the cost of long mixing times. In practice, one must run chains for a vast number of steps to traverse metastable regions, and each transition incurs an expensive energy evaluation, limiting applicability to large-scale systems [32].
Deep generative models, particularly diffusion and flow models tied to continuous-time processes [35, 25], offer a compelling alternative: after training, a cheap constant-cost network evaluation can produce independent samples. Once trained, they do not require energy queries for inference. These models show great success in vision [34], language [24], and audio [21], and can be transferred to unseen conditions. However, their success critically depends on data-driven training on a large number of true equilibrium samples, which are unavailable when only an energy function is known.
In response, researchers have designed new energy-driven algorithms, coined diffusion samplers, to train diffusion-based generative models to sample from the Boltzmann distribution. For example, path integral samplers (PIS) [41] and denoising diffusion samplers (DDS) [37] minimize the divergence between a forward SDE and the backward SDE defined by a Brownian bridge. Akhound-Sadegh et al. [1] proposed iterated denoising energy matching (iDEM), a simulation-free approach to train diffusion models, highlighting the expensive simulation procedure for acquiring new samples from the diffusion-based models. Generative flow networks (GFNs) [3, 22], originating from reinforcement learning, can also be interpreted as a continuous-time Markov process (CTMP) in the limit [6].
However, there still exists a gap between the data- and energy-driven training schemes, particularly on the choice of state spaces and CTMP. Especially, Holderrieth et al. [19] recently proposed generator matching (GM), which allows unified data-driven training of continuous-time processes ranging from flow, diffusion, and jump processes for both continuous and discrete state spaces. This allows training a generative model for hybrid data, e.g. amino acid sequences and 3D coordinates of a protein [7], and greatly expands the available space of parameterization.
@@
[…]
Contribution. We propose energy-based generator matching (EGM), an energy-driven training framework for continuous-time Markov processes parameterized by a neural network. EGM accommodates continuous, discrete, and mixed state spaces and applies to various process types, including diffusion-, flow-, and jump-based models. Our work greatly expands the scope of energy-driven training for continuous-time neural processes (See Table 1).
@@
We validate our work through experiments on various target distributions: discrete Ising model and three joint discrete-continuous tasks, i.e., the Gaussian-Bernoulli restricted Boltzmann machine (GB-RBM) [9], joint double-well potential (JointDW4), and joint mixture of Gaussians (JointMoG). Our findings demonstrate that the training algorithm enables parameterized CTMP to learn the desired distribution and is scalable to reasonably-sized problems.
@@ 2 Preliminary: Generator matching on general state space
In this section, we provide preliminaries on generator matching (GM) [19], which allows generative modeling using arbitrary continuous-time Markov processes (CTMP).
Notation. Let $S$ denote the state space with a reference measure $\nu$. We let $p_{\text{target}}$ denote the target distribution with samples $x_{1}\sim p_{\text{target}}$. We denote a probability measure $p$ on $S$ by $p(dx)$, where “$dx$” indicates integration with respect to $p$ in the variable $x$. If $p$ admits a density, and when no confusion arises, we use $p$ both for the measure $p(dx)$ and its density $p(x):=\frac{dp}{d\nu}(x)$ with respect to $\nu$.
@@ 3 Energy-based generator matching
In this section, we introduce energy-based generator matching (EGM), the first method for learning neural samplers on general state spaces via continuous-time Markov processes. EGM extends the GM framework to energy-driven training by estimating the marginal generator with self-normalized importance sampling (SNIS). To this end, we define a variant of GM loss $L_{\text{GM}}$ in Equation 4:
@@
We first present our scheme to compute the estimator $\hat{F}_{t}$, then describe a bootstrapping technique for variance reduction of the importance weight, and finally provide the training algorithm. Figure 1 provides a visual overview of these estimators. We also include examples demonstrating how EGM accommodates diffusion, flow, and jump processes. [^1] We include the derivation and example in Appendix B for mixed state space.
@@ 3.1 Marginal generator estimation via self-normalized importance sampling
Problem setup. We consider the sampling problem from a density [^2] Though density may not be defined in general, it is given a priori in the sampling problem. $p_{1}(x)=\tilde{p}_{1}(x)/Z$ given only the unnormalized density $\tilde{p}_{1}(x)=\exp(-\mathcal{E}_{1}(x))$ with intractable partition function $Z=\int\tilde{p}_{1}(dx)$. Following GM, our goal is to train a neural network $F_{t}^{\theta}$ to approximate the parametrization $F_{t}$ of the marginal generator $\mathcal{L}_{t}$ in Equation 3.
@@
$\mathcal{L}_{t}f(x)=\mathbb{E}_{x_{1}\sim q_{1|t}(\cdot|x)}\left[w(x,x_{1})\mathcal{L}_{t|1}^{x_{1}}f(x)\right],\quad w(x,x_{1}):=\frac{p_{1|t}(x_{1}|x)}{q_{1|t}(x_{1}|x)}=\frac{\tilde{p}_{1}(x_{1})p_{t|1}(x|x_{1})}{Zp_{t}(x)q_{1|t}(x_{1}|x)}.$ (7)
We avoid computing intractable $Zp_{t}(x)$ for the importance weights by using SNIS scheme:
$\tilde{w}(x,x_{1}):=\frac{\tilde{p}_{1}(x_{1})p_{t|1}(x|x_{1})}{q_{1|t}(x_{1}|x)},\quad w(x,x_{1})=\frac{\tilde{w}(x,x_{1})}{Zp_{t}(x)}=\frac{\tilde{w}(x,x_{1})}{\mathbb{E}_{x_{1}\sim{q_{1|t}(\cdot|x)}}[\tilde{w}(x,x_{1})]}.$ (8)
One can compute the self-normalized weight without knowing the normalization constant $Zp_{t}(x)$. Furthermore, when we use the same set of samples for estimating the generator and normalization constant, we obtain a low-variance estimator at the cost of inducing a bias.
@@
Example 1: Conditional optimal-transport (OT) path. EGM can be applied to train a neural flow sampler driven by the conditional OT path [25]. In the flow model [^3] For a detailed description of flow model and masked diffusion path, see Appendix C., the conditional probability path and its conditional velocity field are expressed by $p_{t|1}(x|x_{1})=\mathcal{N}(x;tx_{1},(1-t)^{2}\text{Id})$ and $u_{t|1}(x|x_{1})=\frac{x_{1}-x}{1-t}$ for $x,x_{1}\in\mathbb{R}^{d}$. Choosing the proposal $q_{1|t}(x_{1}|x)\propto p_{t|1}(x|x_{1})$, one obtains,
@@ 3.2 Bootstrapping tricks for low-variance estimation
The quality of the estimator in Equation 9 depends critically on the proposal distribution $q_{1|t}$. In the ideal case, the proposal $q_{1|t}$ that matches the posterior $p_{1|t}$ ensures that EGM is unbiased and aligns with GM. However, the simple choice of $q_{1|t}(x_{1}|x_{t})\propto p_{t|1}(x_{t}|x_{1})$ leads to a mismatch with the true posterior $p_{1|t}$, resulting in high variance in the importance weight.
@@
Not every $p_{t|r}$ with marginal consistency is tractable. For instance, in the flow model, $p_{t|r}$ is deterministic, making density evaluation infeasible. We give examples of tractable backward transition kernels satisfying our condition for diffusion, flow, and masked-diffusion paths in Appendix C.
Bootstrapped marginalization trick. Now, we generalize Equation 3 and show that the marginal generator $\mathcal{L}_{t}$ can be expressed as marginalization over generators $\mathcal{L}_{t|r}^{x_{r}}$ for probability paths $(p_{t|r}(dx|x_{r}))_{0\leq t\leq r}$ conditioned on time $r$, instead of the conditional generators $\mathcal{L}_{t|1}^{x_{1}}$. To this end, we define marginal consistency of conditional generators as follows:
@@
With generators $\mathcal{L}_{t|r}^{x_{r}}$ conditioned on time $r$ satisfying the consistency, we express the marginal generator $\mathcal{L}_{t}$. We provide the corresponding proof in Appendix A.
Theorem 1
Let $\mathcal{L}_{t|r}^{x_{r}}$ denote the conditional generator for conditional probability path $p_{t|r}(\cdot|x_{r})$ for $0\leq t<r\leq 1$. If the backward transition kernels $p_{t|r}$ satisfy the Equation 12 and the conditional generators satisfy Equation 13, then the marginal generator can be expressed as follows, regardless of $r$:
@@
Similar to Section 3.1, we also consider the estimator based on the self-normalized importance sampling scheme to reduce the variance of the IS estimator. With a simple choice of $q_{r|t}(x_{r}|x_{t})\propto p_{t|r}(x_{t}|x_{r})$, unnormalized importance weight becomes:
@@
^$\hat{\mathcal{L}}_{t}f(x)=\frac{\sum_{i}\tilde{w}(x,x_{r}^{(i)})\mathcal{L}_{t|r}^{x_{r}^{(i)}}f(x)}{\sum_{i}\tilde{w}(x,x_{r}^{(i)})},\qquad\hat{F}_{t}(x)=\frac{\sum_{i}\tilde{w}(x,x_{r}^{(i)})F_{t|r}^{x_{r}^{(i)}}(x)}{\sum_{i}\tilde{w}(x,x_{r}^{(i)})},$ (17)
@@
Intermediate energy estimator. Note that the unnormalized density $\tilde{p}_{r}$ is intractable, unlike $\tilde{p}_{1}$ in Equation 9. Therfore, we train a surrogate $\tilde{p}_{r}^{\phi}(x)$ to approximate the unnormalized density of the intermediate state $x_{r}$. This surrogate can be learned using the following estimator:
[…]
@@
To learn the surrogate, we train on a general version of noised energy matching (NEM) [30] objective:
@@
where $x_{1}^{(1)},\dots,x_{1}^{(K)}$ is sampled from $q_{1|t}$. Then our final loss from Equation 9 becomes:
@@
We check that the transition kernel $p_{t|r}$ satisfies the assumption we proposed in Appendix C.
@@ 3.3 Training details
7:             Update $\phi$ to minimize $L_{\text{NEM}}$ as defined in Equation 20.
@@
9:             Update $\theta$ to minimize $L_{\text{EGM-BS}}$ as defined in Equation 21.
@@
12:             Update $\theta$ to minimize $L_{\text{EGM}}$ as defined in Equation 9.
@@
We now describe the training algorithm for EGM, with the full procedure provided in Algorithm 1.
Bi-level training scheme. To choose a reference distribution $p_{t}^{\text{ref}}$ that is close to $p_{t}$, we use hierarchical sampling procedure $x_{1}\sim\mathcal{B},\,x_{t}\sim p_{t|1}(\cdot|x_{1})$ with replay-buffer $\mathcal{B}$ that approximates $p_{1}$. This leads to the following form of EGM objective:
@@
When the buffer remains close to $p_{1}$, our loss approximates the generator matching loss in Equation 4. We achieve this by continuously updating the buffer, using the bi-level scheme introduced in Akhound-Sadegh et al. [1]. The bi-level scheme alternates between an outer loop and an inner loop, where in the outer loop the buffer $\mathcal{B}$ is improved by drawing samples from the current sampler $F_{t}^{\theta}$. For the inner loop, the sampler $F_{t}^{\theta}$ is trained to minimize the EGM loss with $\mathcal{B}$ held fixed. Because the sampler is updated in the inner loop, the samples collected in the subsequent outer loop reflect its improved performance, thus progressively improving the buffer.
Forward-looking parametrization for masked diffusion. We also introduce a trick to further boost the training of masked diffusion samplers for graphical models via inductive bias on the estimator $\mathcal{E}_{t}^{\phi}$. To this end, consider a graphical model defined on graph $G=(V,E)$ with vertices $V$ and edges $E$. It defines a distribution of vertex-wise variables $x=\{x_{i}\}_{i\in V}$ using an energy function $\mathcal{E}$ expressed as a product of edge-wise potentials $\{\psi_{i,j}\}_{\{i,j\}\in E}$, i.e., $p_{1}(x)\propto\exp\bigl{(}-\mathcal{E}(x)\bigr{)}=\prod_{\{i,j\}\in E}\psi_{i,j}(x^{i},x^{j})$ where we let $x^{i}$ denote the variable associated with vertex $i\in V$.

### v3

@@ 1 Introduction
We tackle the problem of drawing samples from a Boltzmann distribution $p_{\text{target}}(x)\propto\exp(-\mathcal{E}(x))$ given only oracle access to the energy function $\mathcal{E}(x)$ and no pre-computed equilibrium samples. Such “energy-only” inference arises throughout machine learning, e.g., Bayesian inference [gelman1995bayesian] and statistical physics, e.g., computing thermodynamic averages [newman1999monte], yet it remains intractable in high-dimensional or combinatorial state spaces. Classical methods based on Markov processes, e.g., Metropolis–Hastings [metropolis1953equation, hastings1970monte], Gibbs sampling [geman1984stochastic], and Metropolis-adjusted Langevin algorithm [grenander1994representations, roberts1996exponential], provide asymptotically exact guarantees, but at the cost of long mixing times. In practice, one must run chains for a vast number of steps to traverse metastable regions, and each transition incurs an expensive energy evaluation, limiting applicability to large-scale systems [robert1999monte].
Deep generative models, particularly diffusion and flow models tied to continuous-time processes [song2020score, lipmanflow2022], offer a compelling alternative: after training, a cheap constant-cost network evaluation can produce independent samples. Once trained, they do not require energy queries for inference. These models show great success in vision [rombach2022high], language [li2022diffusion], and audio [kong2020diffwave], and can be transferred to unseen conditions. However, their success critically depends on data-driven training on a large number of true equilibrium samples, which are unavailable when only an energy function is known.
In response, researchers have designed new energy-driven algorithms, coined diffusion samplers, to train diffusion-based generative models to sample from the Boltzmann distribution. For example, path integral samplers (PIS) [zhang2021path] and denoising diffusion samplers (DDS) [vargas2023denoising] minimize the divergence between a forward SDE and the backward SDE defined by a Brownian bridge. akhound2024iterated proposed iterated denoising energy matching (iDEM), a simulation-free approach to train diffusion models, highlighting the expensive simulation procedure for acquiring new samples from the diffusion-based models. Generative flow networks (GFNs) [bengio2021flow, lahlou2023theory], originating from reinforcement learning, can also be interpreted as a continuous-time Markov process (CTMP) in the limit [berner2025discrete].
However, there still exists a gap between the data- and energy-driven training schemes, particularly on the choice of state spaces and CTMP. Especially, holderrieth2024generator recently proposed generator matching (GM), which allows unified data-driven training of continuous-time processes ranging from flow, diffusion, and jump processes for both continuous and discrete state spaces. This allows training a generative model for hybrid data, e.g. amino acid sequences and 3D coordinates of a protein [campbell2024generative], and greatly expands the available space of parameterization.
@@
PIS [zhang2021path], iDEM [akhound2024iterated], CMCD [vargas2023transport]
iEFM [woo2024iterated], LFIS [tian2024liouville], NFS2 [chen2025neural]
LEAPS [holderrieth2025leaps]
EGM (ours)
Contribution. We propose energy-based generator matching (EGM), an energy-driven training framework for continuous-time Markov processes parameterized by a neural network. EGM accommodates continuous, discrete, and mixed state spaces and applies to various process types, including diffusion-, flow-, and jump-based models. Our work greatly expands the scope of energy-driven training for continuous-time neural processes (See Table˜1).
@@
We validate our work through experiments on various target distributions: discrete Ising model and three joint discrete-continuous tasks, i.e., the Gaussian-Bernoulli restricted Boltzmann machine (GB-RBM) [cho2011improved], joint double-well potential (JointDW4), and joint mixture of Gaussians (JointMoG). Our findings demonstrate that the training algorithm enables parameterized CTMP to learn the desired distribution and is scalable to reasonably-sized problems.
@@ 2 Preliminary: Generator matching on general state space
In this section, we provide preliminaries on generator matching (GM) [holderrieth2024generator], which allows generative modeling using arbitrary continuous-time Markov processes (CTMP).
Notation. Let $S$ denote the state space with a reference measure $\nu$. We let $p_{\text{target}}$ denote the target distribution with samples $x_{1}\sim p_{\text{target}}$. We denote a probability measure $p$ on $S$ by $p(dx)$, where “dxdx” indicates integration with respect to $p$ in the variable $x$. If $p$ admits a density, and when no confusion arises, we use $p$ both for the measure $p(dx)$ and its density $p(x):=\frac{dp}{d\nu}(x)$ with respect to $\nu$.
@@ 3 Energy-based generator matching
In this section, we introduce energy-based generator matching (EGM), the first method for learning neural samplers on general state spaces via continuous-time Markov processes. EGM extends the GM framework to energy-driven training by estimating the marginal generator with self-normalized importance sampling (SNIS). To this end, we define a variant of GM loss $L_{\text{GM}}$ in Equation˜4:
@@
We first present our scheme to compute the estimator $\hat{F}_{t}$, then describe a bootstrapping technique for variance reduction of the importance weight, and finally provide the training algorithm. Figure˜1 provides a visual overview of these estimators. We also include examples demonstrating how EGM accommodates diffusion, flow, and jump processes. [^1] We include the derivation and example in Appendix B for mixed state space.
@@ 3.1 Marginal generator estimation via self-normalized importance sampling
Problem setup. We consider the sampling problem from a density [^2] Though density may not be defined in general, it is given a priori in the sampling problem. $p_{1}(x)=\tilde{p}_{1}(x)/Z$ given only the unnormalized density $\tilde{p}_{1}(x)=\exp(-\mathcal{E}_{1}(x))$ with intractable partition function $Z=\int\tilde{p}_{1}(dx)$. Following GM, our goal is to train a neural network $F_{t}^{\theta}$ to approximate the parametrization $F_{t}$ of the marginal generator $\mathcal{L}_{t}$ in Equation˜3.
@@
$\mathcal{L}_{t}f(x)=\mathbb{E}_{x_{1}\sim q_{1|t}(\cdot|x)}\left[w(x,x_{1})\mathcal{L}_{t|1}^{x_{1}}f(x)\right],\quad w(x,x_{1}):=\frac{p_{1|t}(x_{1}|x)}{q_{1|t}(x_{1}|x)}=\frac{\tilde{p}_{1}(x_{1})p_{t|1}(x|x_{1})}{Zp_{t}(x)q_{1|t}(x_{1}|x)},$ (7)
where $p_{1|t}$ is the target posterior and $q_{1|t}$ is a proposal kernel. We avoid computing intractable $Zp_{t}(x)$ for the importance weights by using SNIS scheme:
$\tilde{w}(x,x_{1}):=\frac{\tilde{p}_{1}(x_{1})p_{t|1}(x|x_{1})}{q_{1|t}(x_{1}|x)},\quad w(x,x_{1})=\frac{\tilde{w}(x,x_{1})}{Zp_{t}(x)}=\frac{\tilde{w}(x,x_{1})}{\mathbb{E}_{x_{1}\sim{q_{1|t}(\cdot|x)}}[\tilde{w}(x,x_{1})]},$ (8)
where $\tilde{w}$ is the unnormalized importance weight that is tractable to compute. In this way, one can compute the self-normalized weight without knowing the normalization constant $Zp_{t}(x)$. Furthermore, when we use the same samples for estimating the generator and normalization constant, we obtain a low-variance estimator at the cost of inducing a bias.
@@
Example 1: Conditional optimal-transport (OT) path. EGM can be applied to train a neural flow sampler driven by the conditional OT path [lipmanflow2022]. In the flow model [^3] For a detailed description of flow model and masked diffusion path, see Appendix C., the conditional probability path and its conditional velocity field are expressed by $p_{t|1}(x|x_{1})=\mathcal{N}(x;tx_{1},(1-t)^{2}\text{Id})$ and $u_{t|1}(x|x_{1})=\frac{x_{1}-x}{1-t}$ for $x,x_{1}\in\mathbb{R}^{d}$. Choosing the proposal $q_{1|t}(x_{1}|x)\propto p_{t|1}(x|x_{1})$, one obtains,
@@ 3.2 Bootstrapping tricks for low-variance estimation
The quality of the estimator in Equation˜9 depends critically on the proposal distribution $q_{1|t}$. In the ideal case, the proposal $q_{1|t}$ that matches the posterior $p_{1|t}$ ensures that EGM is unbiased and aligns with GM. However, the simple choice of $q_{1|t}(x_{1}|x_{t})\propto p_{t|1}(x_{t}|x_{1})$ leads to a mismatch with the true posterior $p_{1|t}$, resulting in high variance in the importance weight.
@@
Not every $p_{t|r}$ with marginal consistency is tractable. For instance, in the flow model, $p_{t|r}$ is deterministic, making density evaluation infeasible. We give examples of tractable backward transition kernels satisfying our condition for diffusion, flow, and masked-diffusion paths in Appendix˜C.
Bootstrapped marginalization trick. Now, we generalize Equation˜3 and show that the marginal generator $\mathcal{L}_{t}$ can be expressed as marginalization over generators $\mathcal{L}_{t|r}^{x_{r}}$ for probability paths $(p_{t|r}(dx|x_{r}))_{0\leq t\leq r}$ conditioned on time $r$, instead of the conditional generators $\mathcal{L}_{t|1}^{x_{1}}$. To this end, we define marginal consistency of conditional generators as follows:
@@
With generators $\mathcal{L}_{t|r}^{x_{r}}$ conditioned on time $r$ satisfying the consistency, we express the marginal generator $\mathcal{L}_{t}$. We provide the corresponding proof in Appendix˜A.
Theorem 1.
Let $\mathcal{L}_{t|r}^{x_{r}}$ denote the conditional generator for conditional probability path $p_{t|r}(\cdot|x_{r})$ for $0\leq t<r\leq 1$. If the backward transition kernels $p_{t|r}$ satisfy the Equation˜12 and the conditional generators satisfy Equation˜13, then the marginal generator can be expressed as follows, regardless of $r$:
@@
Similar to Section˜3.1, we also consider the estimator based on the self-normalized importance sampling scheme to reduce the variance of the IS estimator. With a simple choice of $q_{r|t}(x_{r}|x_{t})\propto p_{t|r}(x_{t}|x_{r})$, unnormalized importance weight becomes:
@@
$\hat{\mathcal{L}}_{t}f(x)=\frac{\sum_{i}\tilde{w}(x,x_{r}^{(i)})\mathcal{L}_{t|r}^{x_{r}^{(i)}}f(x)}{\sum_{i}\tilde{w}(x,x_{r}^{(i)})},\qquad\hat{F}_{t}(x)=\frac{\sum_{i}\tilde{w}(x,x_{r}^{(i)})F_{t|r}^{x_{r}^{(i)}}(x)}{\sum_{i}\tilde{w}(x,x_{r}^{(i)})},$ (17)
@@
Intermediate energy estimator. Note that the unnormalized density $\tilde{p}_{r}$ is intractable, unlike $\tilde{p}_{1}$ in Equation˜9. Therfore, we train a surrogate $\tilde{p}_{r}^{\phi}(x)$ to approximate the unnormalized density of the intermediate state $x_{r}$. This surrogate can be learned using the following estimator:
$\tilde{p}_{r}(x_{r}):=\int p_{r|1}(x_{r}|x_{1})\tilde{p}_{1}(x_{1})dx_{1}=Z_{1|r}(x_{r})\mathbb{E}_{x_{1}\sim q_{1|r}(\cdot|x_{r})}[\tilde{p}_{1}(x_{1})],$ (18)
@@
To learn the surrogate, we train on a general version of noised energy matching (NEM) [ouyang2024bnem] objective:
@@
where $x_{1}^{(1)},\dots,x_{1}^{(K)}$ is sampled from $q_{1|t}$. Then our final loss from Equation˜9 becomes:
@@
We check that the transition kernel $p_{t|r}$ satisfies the assumption we proposed in Appendix˜C.
@@ 3.3 Training details
7:     Update $\phi$ to minimize $L_{\text{NEM}}$ as defined in Equation˜20.
@@
9:     Update $\theta$ to minimize $L_{\text{EGM-BS}}$ as defined in Equation˜21.
@@
12:     Update $\theta$ to minimize $L_{\text{EGM}}$ as defined in Equation˜9.
@@
We now describe the training algorithm for EGM, with the full procedure provided in Algorithm˜1.
Bi‑level training scheme. To choose a reference distribution $p_{t}^{\text{ref}}$ that is close to $p_{t}$, we use hierarchical sampling procedure $x_{1}\sim\mathcal{B},\,x_{t}\sim p_{t|1}(\cdot|x_{1})$ with replay-buffer $\mathcal{B}$ that approximates $p_{1}$. This leads to the following form of EGM objective:
@@
When the buffer remains close to $p_{1}$, our loss approximates the generator matching loss in Equation˜4. We achieve this by continuously updating the buffer, using the bi-level scheme introduced in akhound2024iterated. The bi-level scheme alternates between an outer loop and an inner loop, where in the outer loop the buffer $\mathcal{B}$ is improved by drawing samples from the current sampler $F_{t}^{\theta}$. For the inner loop, the sampler $F_{t}^{\theta}$ is trained to minimize the EGM loss with $\mathcal{B}$ held fixed. Because the sampler is updated in the inner loop, the samples collected in the subsequent outer loop reflect its improved performance, thus progressively improving the buffer.
Forward‑looking parametrization for masked diffusion. We also introduce a trick to further boost the training of masked diffusion samplers for graphical models via inductive bias on the estimator $\mathcal{E}_{t}^{\phi}$. To this end, consider a graphical model defined on graph $G=(V,E)$ with vertices $V$ and edges $E$. It defines a distribution of vertex-wise variables $x=\{x_{i}\}_{i\in V}$ using an energy function $\mathcal{E}$ expressed as a product of edge-wise potentials $\{\psi_{i,j}\}_{\{i,j\}\in E}$, i.e., $p_{1}(x)\propto\exp\bigl(-\mathcal{E}(x)\bigr)=\prod_{\{i,j\}\in E}\psi_{i,j}(x^{i},x^{j})$ where we let $x^{i}$ denote the variable associated with vertex $i\in V$.